# Projective limits of probabilistic symmetries and their applications to random graph limits

### v1

@@ 4.1 Random graphs as point processes
Conversely, any (random) symmetric measure $\xi=\sum_{i}\delta_{e_{i}}$, where $e_{i}\in\mathbb{X}$, has an associated (random) undirected graph with the edge set $E=\cup_{i}e_{i}$. Therefore, we will identify the space of symmetric measures $M_{s}(\mathbb{X})$ on space $\mathbb{X}=\mathbb{L}\times\mathbb{L}$ with the space of undirected graphs with vertex labels in $\mathbb{L}$. Technically, in order to have $\xi^{G}$ be locally finite we need that for any bounded set $B\subseteq\mathbb{L}$ the set of vertices in $B$ is finite, i.e., $|B\cap V|<\infty$. This will be true for all the examples considered in this section, so we choose to skip the additional technical definitions here.
@@ 4.3 Specific examples of projective graph limits
      the invariance with respect to $\Phi_{\infty}$ is of finite characteristic.
@@ 4.3.1 Graphons as projective limits of random graphs on $\mathbb{N}$, invariant under permutations
      We let $\mathbb{L}=\mathbb{N}$ and $\mathbb{L}_{n}=[n]:=\{1,2,\dots,n\}$, and endow each of these with the standard counting topology.
@@
      Since we deal with with the standard counting topology, $\mathbb{L}$ is Polish as required. Moreover, we observe that each $\mathbb{L}_{n}$ is indeed an open subset of $\mathbb{L}$, and clearly $(\mathbb{L}_{n})_{n\geq 1}$ is an exhausting non-decreasing sequence.
@@
      It is known [21] that a measure on $\mathbb{N}\times\mathbb{N}$ is invariant under permutations of $\mathbb{N}$ if it is invariant under any permutation $\varphi_{n}\in\Phi_{n}$ for all $n\geq 1$, so $\Psi_{n}=\Phi_{n}$ in this case.
Therefore, the application of Corollaries 4.1 and 4.2 to these specific settings results in that the projective limit of any projective system of random graphs on $[n]$ that are invariant under finite permutations is a random graph on $\mathbb{N}$ that is invariant under infinite permutations. Furthermore, by evoking the classical representation result by Aldous and Hoover [3, 26], this limit random graph is a probabilistic mixture of $W$-random graphs, where $W:[0,1]\times[0,1]\to[0,1]$ is a symmetric measurable function known as a graphon. A $W$-random graph is constructed by considering a sequence $x_{1},x_{2},\dots$ of i.i.d. points sampled from $U[0,1]$, and connecting vertices $i$ and $j$ independently with probability $W(x_{i},x_{j})$.
@@
      For any $j,k\in\mathbb{N}$, let $I_{j,k}$ be the $j$’th $k$-dyadic interval $I_{j,k}=((j-1)/2^{k},j/2^{k}]$. For any two dyadic intervals $I_{j,k},I_{i,k}\subset\mathbb{R}_{+}$, denote by $\vartheta_{i,j,k}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ the function that swaps them. Observe that any $\vartheta_{i,j,k}$ is a measure-preserving transformation of $[0,m)$ with $m=\max\{i,j\}/2^{k}$, and hence it is an element of $\Phi_{m}$. Denote by $\Psi_{1}$ the set of swaps of dyadic intervals contained in $[0,1)$, and given $\Psi_{n}$, let $\Psi_{n+1}$ be all swaps of dyadic intervals contained in $[0,n+1)$ that are not in $\Psi_{n}$. Observe that the sequence $\Psi_{n}$ provides the finite characteristic for measure-preserving transformations of $\mathbb{R}_{+}$ thanks to the fact, Proposition 9.1 in [30], that a random measure on $\mathbb{R}_{+}\times\mathbb{R}_{+}$ is invariant under measure-preserving transformations of $\mathbb{R}_{+}$ if and only if it is invariant under $\vartheta_{i,j,k}$ for all $i,j,k\in\mathbb{N}$.
Therefore, the application of Corollaries 4.1 and 4.2 to these specific settings results in that the projective limit of any projective system of random graphs on $[0,n)$ that are invariant under measure-preserving transformations is a random graph on $\mathbb{R}_{+}$ that is invariant under measure-preserving transformations. Furthermore, by evoking the representation result [11, Theorem 2.8] or [47, Theorem 4.7]), which is a direct consequence of a general representation result for jointly exchangeable random measures on $\mathbb{R}_{+}\times\mathbb{R}_{+}$ by Kallenberg [30], this limit random graph is a probabilistic mixture of $\mathcal{W}$-random graphs, where $\mathcal{W}$ is a graphex. A graphex is a triple $\mathcal{W}=(I,S,W)$, where $I\in\mathbb{R}_{+}$, $S:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is measurable with $S\wedge 1$ being integrable, and $W:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a symmetric measurable function with some additional conditions that we do not go into for brevity. To sample a $\mathcal{W}$-random graph on $[0,n)$, one samples the unit-rate Poisson point process $(x_{i},y_{i})$ on $0,n)\times\mathbb{R}_{+}$, creates edges between vertices labeled by $x_{i}$ and $x_{j}$ independently with probability $W(y_{i},y_{j})$, and then removes zero-degree vertices. The $I$ and $S$ components of the graphex can be ignored. See the references above for further details.
@@ 5 Proofs
were we used compatibility in the form of diagram (6) for the third equality.
@@
The main object of interest concerns the space $M(\mathbb{X})$ of locally finite simple (symmetric) counting measures on $\mathbb{X}$, equipped with the topology $\mathcal{T}$ generated by the evaluation maps $\xi\mapsto\xi(B)$. We will denote by $\mathcal{T}_{n}$ the subspace topology for the subspace $M(\mathbb{X}_{n})$.
The proof of Theorem 3.2 consists of two main steps. First we use a general approach to construct the projective limit of the system $\langle M(\mathbb{X}_{n}),\mathcal{T}_{n},\pi_{mn}\rangle_{\mathbb{N}}$ and prove that this space is isomorphic to $(M(\mathbb{X}),\mathcal{T})$.
For the second step, we use a known result which guarantees the a projective sequence of probability measures $\mu_{n}$ on $(M(\mathbb{X}_{n}),)$ has a projective limit. Since the projective limit space is isomorphic to $(M(\mathbb{X}),\mathcal{T})$ we then get a probability measure on this space, which establishes the result.
We first show that in our setting the projective limit is isomorphic to $(M(\mathbb{X}),\mathcal{T})$. Recall that we defined the projections $\pi_{n}:M(\mathbb{X})\to M(\mathbb{X}_{n})$ via the canonical maps $\iota_{n}:\mathbb{X}_{n}\to\mathbb{X}$ as
$\pi_{n}\xi=\xi\circ\iota_{n}.$
@@
Let $\langle M(\mathbb{X}_{n}),\mathcal{T}_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of locally finite simple (symmetric) counting measures, with $\pi_{mn}$ defined as in (9), and let $\langle M_{\mathbb{N}},\mathcal{T}_{\mathbb{N}},\phi_{n}\rangle$ be its projective limit as defined in Theorem 2.1. Then there exist an isomorphism of topological spaces
$h:(M_{\mathbb{N}},\mathcal{T}_{\mathbb{N}})\to(M(\mathbb{X}),\mathcal{M}).$
@@
First we need to prove that $h(\vec{\xi})$ is indeed a measure. We start by observing that it satisfies non-negativity and $h(\vec{\xi})(\emptyset)=0$. In addition, it is also $\sigma$-additive. For this consider a sequence $(A_{k})_{k\geq 1}$ in $\mathcal{X}$ such that $A_{k+1}\subset A_{k}$ and $A_{k}\downarrow\emptyset$ and set $A_{k,n}^{\mathbb{X}}=A_{k}\cap\mathbb{X}_{n}$. Then there exists a $K_{0}$ such that for all $k\geq K_{0}$, $A_{k}\subseteq\mathbb{X}_{1}$ and thus $A_{k,n}^{\mathbb{X}}=A_{k,1}^{\mathbb{X}}$ for all $k\geq K_{0}$ and $n\geq 1$.
But then, since $\xi_{n}(A)=\xi_{1}(A)$ for all $A\in\mathcal{X}_{1}$, this implies that
$\lim_{k\to\infty}h(\vec{\xi})(A_{k})=\lim_{k\to\infty}\lim_{n\to\infty}\xi_{n}(A_{k,n}^{\mathbb{X}})=\lim_{k\to\infty}\lim_{n\to\infty}\xi_{n}(A_{k,1}^{\mathbb{X}})=\lim_{k\to\infty}\xi_{1}(A_{k,1}^{\mathbb{X}})=0.$
Hence, $h(\vec{\xi})$ is $\sigma$-additive. In particular we conclude that $h(\vec{\xi})$ is a measure on $\mathbb{X}$. Finally, we need to show that $h(\vec{\xi})$ is locally finite. For this let $A\in\mathcal{X}$ be a bounded measurable set. Then, there must exist a $N\in\mathbb{N}$ such that $A\subset\mathbb{X}_{N}$ and hence $h(\vec{\xi})(A)=\xi_{N}(A)<\infty$ since each $\xi_{n}$ was a locally finite measure. We thus conclude that $h(\vec{\xi})\in M(\mathbb{X})$.
@@
Consider a sequence $(\vec{\xi}^{k})_{k\geq 1}$ of elements of $M_{\mathbb{N}}$ such that $\vec{\xi}^{k}\to\vec{\xi}^{\infty}$. It then suffices to show that $h(\vec{\xi}^{k})\to h(\vec{\xi}^{\infty})$ in $M(\mathbb{X})$. To this end let $B$ be a measurable set of $\mathbb{X}$ and note that by definition of the topology $\mathcal{T}_{\mathbb{N}}$ is holds that for each $n\geq 1$, $\xi^{k}_{n}(B\cap\mathbb{X}_{n})\to\xi_{n}^{\infty}(B\cap\mathbb{X}_{n})$ as $k\to\infty$. ,
Fix $\varepsilon>0$ and $N\geq 1$. Then there exists a $K_{N}$ such that for all $k\geq K_{N}$,
$\xi_{N}^{\infty}(B\cap\mathbb{X}_{N})-\xi_{N}^{k}(B\cap\mathbb{X}_{N})<\varepsilon.$
Moreover, since the sequence $(\xi^{k}_{n}(B\cap\mathbb{X}_{n}))_{n\geq 1}$ is non-decreasing for any $k$, it follows that for all $k>K_{N}$ and $n\geq N$
$\xi_{N}^{\infty}(B\cap\mathbb{X}_{N})-\xi_{n}^{k}(B\cap\mathbb{X}_{n})\leq\xi_{N}^{\infty}(B\cap\mathbb{X}_{N})-\xi_{N}^{k}(B\cap\mathbb{X}_{N})<\varepsilon.$
Hence we conclude that
$\lim_{k\to\infty}h(\vec{\xi}_{k})(B)$ $=\lim_{k\to\infty}\lim_{n\to\infty}\xi_{n}^{k}(B\cap\mathbb{X}_{n})$
$=\lim_{n\to\infty}\lim_{k\to\infty}\xi_{n}^{k}(B\cap\mathbb{X}_{n})$
$=\lim_{n\to\infty}\xi_{n}^{\infty}(B\cap\mathbb{X}_{n})=h(\vec{\xi}^{\infty})(B),$
which implies that $h(\vec{\xi}^{k})\to h(\vec{\xi}^{\infty})$ in $M(\mathbb{X})$ and thus proves that $h$ is continuous.
@@
It now remains to show that $g$ is continuous. Let $A\in\mathcal{T}:=\bigcup_{n\geq 1}\phi_{n}^{-1}(\mathcal{T}_{n})$ so that $A=\phi_{n}^{-1}(A_{n})$ for some $A_{n}\in\mathcal{T}_{n}$. Then, noting that the commuting diagram implies that $\phi_{n}g=\pi_{n}$, we have
$g^{-1}(A)=g^{-1}(\phi_{n}^{-1}(A_{n}))=(\phi_{n}g)^{-1}(A_{n})=\pi_{n}^{-1}(A_{n})\in T,$
so that $g$ is continuous.
@@
Let use denote by $\mathcal{M}_{n}$ the subspace $\sigma$-algebra of $M(\mathbb{X}_{n})$. Then by Theorem 2.2 any projective sequence $(\mu_{n})_{n\geq 1}$ of probability measures on $(M(\mathbb{X}_{n}),\mathcal{M}_{n})$ has a projective limit $\hat{\mu}_{\infty}$ which is a probability measure on the projective limit space $(M_{\mathbb{N}},\sigma(\mathcal{T}_{\mathbb{N}}))$, defined as in Theorem 2.1. By Proposition 5.1 this space is isomorphic to $(M(\mathbb{X}),\mathcal{T})$ via the isomorphism $h$ and hence the probability measure $\hat{\mu}_{\infty}$ can be extended by means of a push forward to a probability measure $\mu_{\mathbb{N}}:=h\ast\hat{\mu}_{\infty}$ on this space.
@@
We need to show that the projective limit $\mu_{\mathbb{N}}$ is invariant with respect to the compatible direct pre-limit $\Phi_{\infty}$. Recall that we have extended the action of $\Phi_{n}$ from $\mathbb{X}_{n}$ to $M(\mathbb{X}_{n})$ by defining $\varphi_{n}\xi_{n}(A)=\xi_{n}(\varphi_{n}^{-1}(A))$ and similarly for the action of $\Phi_{\infty}$ on $M(\mathbb{X})$.
Because the invariance has a finite characteristic, it suffice to show that $\mu_{\mathbb{N}}$ is invariant with respect to $\eta_{m}\varphi_{m}$ for any $m\geq 1$ and $\varphi_{m}\in\Phi_{m}$. For this we follow a strategy similar to the proof of invariance in Theorem 3.1, making use of the fact that the projective limit measure $\mu_{\mathbb{N}}$ is the unique probability measure such that $\pi_{n}\ast\mu_{\mathbb{N}}=\mu_{n}$.
@@
Fix a $m\geq 1$, $\varphi_{m}\in\Psi_{m}\subseteq\Phi_{m}$ and define $\hat{\mu}_{\infty}=\mu_{\mathbb{N}}\circ(\eta_{m}\varphi_{m})^{-1}$. Then we need to show that $\hat{\mu}_{\infty}=\mu_{\mathbb{N}}$, and by the uniqueness of projective limits it suffices to show that $\pi_{n}*\hat{\mu}_{\infty}=\mu_{n}$ holds for all $n\geq 1$. Let $k=\max\{m,n\}$ and $A_{n}\in\mathcal{M}_{n}$ be any measurable set of graphs with node labels in $\mathbb{X}_{n}$. Then,
@@
    were we used the commuting diagram in point 2 of the lemma for the third step.

### v2

@@ 4.1 Random graphs as point processes
The isolated vertices labeled by $z\in\mathbb{L}$ in $G$ are represented by self-loop edge-points $\delta_{(z,z)}$. Conversely, any (random) symmetric measure $\xi=\sum_{i}\delta_{e_{i}}$, where $e_{i}\in\mathbb{X}$, has an associated (random) undirected graph with the edge set $E=\cup_{i}e_{i}$ and vertex set $V=\{x\in\mathbb{L}\,:\,\exists i\text{ and }y\in\mathbb{L}\text{ such that }e_{i}=(x,y)\}$. Therefore, we will identify the space of symmetric measures $M_{s}(\mathbb{X})$ on space $\mathbb{X}=\mathbb{L}\times\mathbb{L}$ with the space of undirected graphs and vertex labels in $\mathbb{L}$. Technically, in order to have $\xi^{G}$ be locally finite we need that for any bounded set $B\subseteq\mathbb{L}$ the set of vertices in $B$ is finite, i.e., $|B\cap V|<\infty$. This will be true for all the examples considered in this section, so we choose to skip the additional technical definitions here.
@@ 4.3 Specific examples of projective graph limits
      invariance with respect to each $\Phi_{n}$ implies invariance with respect to $\Phi_{\infty}$.
@@ 4.3.1 Graphons as projective limits of random graphs on $\mathbb{N}$, invariant under permutations
      We let $\mathbb{L}=\mathbb{N}$ and $\mathbb{L}_{n}=[n]:=\{1,2,\dots,n\}$, and endow each of these with the standard discrete topology.
@@
      Since we deal with the standard discrete topology, $\mathbb{L}$ is Polish as required. Moreover, we observe that each $\mathbb{L}_{n}$ is indeed an open subset of $\mathbb{L}$, and clearly $(\mathbb{L}_{n})_{n\geq 1}$ is an exhausting non-decreasing sequence.
@@
      It is known [21] that a measure on $\mathbb{N}\times\mathbb{N}$ is invariant under permutations of $\mathbb{N}$ if it is invariant under any permutation $\varphi_{n}\in\Phi_{n}$ for all $n\geq 1$.
Therefore, the application of Corollaries 4.1 and 4.2 to these specific settings results in that the projective limit of any projective system of random graphs on $[n]$ that are invariant under finite permutations is a random graph on $\mathbb{N}$ that is invariant under infinite permutations. Furthermore, by invoking the classical representation result by Aldous and Hoover [3, 26], this limit random graph is a probabilistic mixture of $W$-random graphs, where $W:[0,1]\times[0,1]\to[0,1]$ is a symmetric measurable function known as a graphon. A $W$-random graph is constructed by considering a sequence $x_{1},x_{2},\dots$ of i.i.d. points sampled from $U[0,1]$, and connecting vertices $i$ and $j$ independently with probability $W(x_{i},x_{j})$.
@@
      A random measure on $\mathbb{R}_{+}\times\mathbb{R}_{+}$ is invariant under measure-preserving transformations of $\mathbb{R}_{+}$ if and only if it is invariant under transpositions of dyadic intervals, see Proposition 9.1 in [30]. Since any two such intervals are finite and thus contained in $[0,m]$ for some $m$, it follows that invariance with respect to $\Phi_{n}$ for all $n\geq 1$ implies invariance with respect to $\Phi_{\infty}$.
Therefore, the application of Corollaries 4.1 and 4.2 to these specific settings results in that the projective limit of any projective system of random graphs on $[0,n)$ that are invariant under measure-preserving transformations is a random graph on $\mathbb{R}_{+}$ that is invariant under measure-preserving transformations. Furthermore, by invoking the representation result [11, Theorem 2.8] or [48, Theorem 4.7]), which is a direct consequence of a general representation result for jointly exchangeable random measures on $\mathbb{R}_{+}\times\mathbb{R}_{+}$ by Kallenberg [30], this limit random graph is a probabilistic mixture of $\mathcal{W}$-random graphs, where $\mathcal{W}$ is a graphex. A graphex is a triple $\mathcal{W}=(I,S,W)$, where $I\in\mathbb{R}_{+}$, $S:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is measurable with $S\wedge 1$ being integrable, and $W:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a symmetric measurable function with some additional conditions that we do not go into for brevity. To sample a $\mathcal{W}$-random graph on $[0,n)$, one samples the unit-rate Poisson point process $(x_{i},y_{i})$ on $0,n)\times\mathbb{R}_{+}$, creates edges between vertices labeled by $x_{i}$ and $x_{j}$ independently with probability $W(y_{i},y_{j})$, and then removes zero-degree vertices. The $I$ and $S$ components of the graphex can be ignored. See the references above for further details.
@@ 5 Proofs
where we used compatibility in the form of diagram (6) for the third equality.
@@
The main object of interest is the space $M(\mathbb{X})$ of locally finite simple (symmetric) counting measures on $\mathbb{X}$, equipped with the vague topology $\mathcal{T}$, i.e., the coarsest topology such that the maps $\xi\longmapsto\int f\,d\xi$, $f\in C_{c}(X)$, are continuous, where $C_{c}(X)$ is the space of continuous functions on $X$ with compact support. We denote by $\mathcal{T}_{n}$ the subspace topology on $M(\mathbb{X}_{n})$ induced by $\mathcal{T}$. Recall that $\mathcal{M}$ is the $\sigma$-algebra on $M(\mathbb{X})$ introduced in Section 2.1 and that this coincides with the Borel $\sigma$-algebra of the vague topology [31].
The proof of Theorem 3.2 consists of two main steps. First, we use the general constructions from Section 2.2.1 to build the projective limit of the system $\langle M(\mathbb{X}_{n}),\mathcal{T}_{n},\pi_{mn}\rangle_{\mathbb{N}}$ and prove that this space is homeomorphic to $(M(\mathbb{X}),\mathcal{T})$.
For the second step, we apply Theorem 2.2 to obtain a projective limit measure for any projective sequence of probability measures $\mu_{n}$ on $(M(\mathbb{X}_{n}),\mathcal{M}_{n})$, where $\mathcal{M}_{n}$ is the subspace $\sigma$-algebra on $M(\mathbb{X}_{n})$. Since the projective limit space is homeomorphic to $(M(\mathbb{X}),\mathcal{T})$, we then get a probability measure on this space, which establishes the result.
We first show that in our setting the projective limit is homeomorphic to $(M(\mathbb{X}),\mathcal{T})$. Recall that we defined the projections $\pi_{n}:M(\mathbb{X})\to M(\mathbb{X}_{n})$ via the canonical maps $\iota_{n}:\mathbb{X}_{n}\to\mathbb{X}$ as
$\pi_{n}\xi(A_{n})=\xi(\iota_{n}(A_{n})).$
@@
Let $\langle M(\mathbb{X}_{n}),\mathcal{T}_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of locally finite simple (symmetric) counting measures, with $\pi_{mn}$ defined as in (9), and let $\langle M_{\mathbb{N}},\mathcal{T}_{\mathbb{N}},\phi_{n}\rangle$ be its projective limit as defined in Theorem 2.1. Then there exists an isomorphism of topological spaces
$h:(M_{\mathbb{N}},\mathcal{T}_{\mathbb{N}})\to(M(\mathbb{X}),\mathcal{T}).$
@@
First we need to prove that $h(\vec{\xi})$ is indeed a measure. For each $n\geq 1$ define a measure $\bar{\xi}_{n}$ on $(\mathbb{X},\mathcal{X})$ by
$\bar{\xi}_{n}(A):=\xi_{n}(A\cap\mathbb{X}_{n}),\qquad A\in\mathcal{X}.$
Clearly each $\bar{\xi}_{n}$ is a (simple counting) measure on $\mathbb{X}$. Moreover, for $m\geq n$ and $A\in\mathcal{X}$ we have, using $\mathbb{X}_{n}\subseteq\mathbb{X}_{m}$ and the projective consistency $\xi_{n}=\pi_{mn}\xi_{m}$,
$\bar{\xi}_{n}(A)=\xi_{n}(A\cap\mathbb{X}_{n})=\pi_{mn}\xi_{m}(A\cap\mathbb{X}_{n})=\xi_{m}(\iota_{nm}(A\cap\mathbb{X}_{n}))\leq\xi_{m}(A\cap\mathbb{X}_{m})=\bar{\xi}_{m}(A).$
Hence $(\bar{\xi}_{n}(A))_{n\geq 1}$ is non-decreasing for every $A\in\mathcal{X}$, and by definition
$h(\vec{\xi})(A)=\lim_{n\to\infty}\bar{\xi}_{n}(A),\qquad A\in\mathcal{X}.$
We now verify that $h(\vec{\xi})$ is a measure. Let $(A_{i})_{i\geq 1}\subseteq\mathcal{X}$ be pairwise disjoint. For each $n$ we have $\bar{\xi}_{n}\!\left(\bigcup_{i\geq 1}A_{i}\right)=\sum_{i\geq 1}\bar{\xi}_{n}(A_{i})$, and therefore
$h(\vec{\xi})\!\left(\bigcup_{i\geq 1}A_{i}\right)=\lim_{n\to\infty}\bar{\xi}_{n}\!\left(\bigcup_{i\geq 1}A_{i}\right)=\lim_{n\to\infty}\sum_{i\geq 1}\bar{\xi}_{n}(A_{i}).$
Let $S_{n,M}:=\sum_{i=1}^{M}\bar{\xi}_{n}(A_{i})$. Then $S_{n,M}$ is non-decreasing in both $n$ and $M$, and $\sum_{i\geq 1}\bar{\xi}_{n}(A_{i})=\sup_{M\geq 1}S_{n,M}$. Using monotonicity in both indices we obtain
$\lim_{n\to\infty}\sum_{i\geq 1}\bar{\xi}_{n}(A_{i})=\lim_{n\to\infty}\sup_{M\geq 1}S_{n,M}=\sup_{M\geq 1}\lim_{n\to\infty}S_{n,M}=\sup_{M\geq 1}\sum_{i=1}^{M}\lim_{n\to\infty}\bar{\xi}_{n}(A_{i})=\sum_{i\geq 1}h(\vec{\xi})(A_{i}),$
so $h(\vec{\xi})$ is countably additive. In particular, $h(\vec{\xi})$ is a measure on $\mathbb{X}$.
Finally, we need to show that $h(\vec{\xi})$ is locally finite. For this let $A\in\mathcal{X}$ be a bounded measurable set. By the exhaustion property of $(\mathbb{X}_{n})_{n\geq 1}$ there exists $N\in\mathbb{N}$ such that $A\subseteq\mathbb{X}_{N}$, and hence
$h(\vec{\xi})(A)=\lim_{n\to\infty}\xi_{n}(A\cap\mathbb{X}_{n})=\xi_{N}(A)<\infty,$
since $\xi_{N}$ is locally finite. We thus conclude that $h(\vec{\xi})\in M(\mathbb{X})$.
@@
Let $(\vec{\xi}^{k})_{k\geq 1}$ be a sequence in $M_{\mathbb{N}}$ such that $\vec{\xi}^{k}\to\vec{\xi}^{\infty}$ in $(M_{\mathbb{N}},\mathcal{T}_{\mathbb{N}})$. By definition of the projective limit topology $\mathcal{T}_{\mathbb{N}}$ (the initial topology w.r.t. the maps $\phi_{n}$), we have, for every fixed $n\geq 1$,
$\xi_{n}^{k}=\phi_{n}(\vec{\xi}^{k})\ \longrightarrow\ \phi_{n}(\vec{\xi}^{\infty})=\xi_{n}^{\infty}\qquad\text{in }(M(\mathbb{X}_{n}),\mathcal{T}_{n}).$
To prove that $h(\vec{\xi}^{k})\to h(\vec{\xi}^{\infty})$ in $(M(\mathbb{X}),\mathcal{T})$ with $\mathcal{T}$ the vague topology, it suffices to show that
$\int_{\mathbb{X}}f\,d\,h(\vec{\xi}^{k})\ \longrightarrow\ \int_{\mathbb{X}}f\,d\,h(\vec{\xi}^{\infty})\qquad\text{for all }f\in C_{c}(\mathbb{X}).$
Fix $f\in C_{c}(\mathbb{X})$. Since $\mathrm{supp}(f)$ is compact and $(\mathbb{X}_{n})_{n\geq 1}$ exhausts $\mathbb{X}$, there exists $N$ such that $\mathrm{supp}(f)\subseteq\mathbb{X}_{N}$. For any $\vec{\xi}\in M_{\mathbb{N}}$, the measure $h(\vec{\xi})$ agrees with $\xi_{N}$ on $\mathbb{X}_{N}$ (indeed, for $A\subseteq\mathbb{X}_{N}$ we have $A\cap\mathbb{X}_{n}=A$ for all $n\geq N$ and $(\pi_{nN}\xi_{n})(A)=\xi_{N}(A)$, hence $\xi_{n}(A)=\xi_{N}(A)$). It follows then that
$\int_{\mathbb{X}}f\,d\,h(\vec{\xi})=\int_{\mathbb{X}_{N}}f\,d\xi_{N}.$
Applying this to $\vec{\xi}^{k}$ and $\vec{\xi}^{\infty}$ gives
$\int_{\mathbb{X}}f\,d\,h(\vec{\xi}^{k})=\int_{\mathbb{X}_{N}}f\,d\xi_{N}^{k},\qquad\int_{\mathbb{X}}f\,d\,h(\vec{\xi}^{\infty})=\int_{\mathbb{X}_{N}}f\,d\xi_{N}^{\infty}.$
Since $\xi_{N}^{k}\to\xi_{N}^{\infty}$ in $(M(\mathbb{X}_{N}),\mathcal{T}_{N})$ and $\mathcal{T}_{N}$ is the vague topology on $M(\mathbb{X}_{N})$, we obtain
$\int_{\mathbb{X}_{N}}f\,d\xi_{N}^{k}\ \longrightarrow\ \int_{\mathbb{X}_{N}}f\,d\xi_{N}^{\infty},$
and therefore $\int f\,d\,h(\vec{\xi}^{k})\to\int f\,d\,h(\vec{\xi}^{\infty})$ for all $f\in C_{c}(\mathbb{X})$. This is exactly $h(\vec{\xi}^{k})\to h(\vec{\xi}^{\infty})$ in the vague topology on $M(\mathbb{X})$, so $h$ is continuous.
@@
It now remains to show that $g$ is continuous. Since $\mathcal{T}_{\mathbb{N}}$ is the initial topology with respect to the maps $\phi_{n}$, it suffices to show that $\phi_{n}\circ g$ is continuous for each $n\geq 1$. By the commuting diagram we have $\phi_{n}g=\pi_{n}$, where $\pi_{n}$ is the restriction map $\pi_{n}(\xi)=\xi|_{\mathbb{X}_{n}}$.
We claim that $\pi_{n}:(M(\mathbb{X}),\mathcal{T})\to(M(\mathbb{X}_{n}),\mathcal{T}_{n})$ is continuous for the vague topologies. Indeed, let $f\in C_{c}(\mathbb{X}_{n})$ and choose $\tilde{f}\in C_{c}(\mathbb{X})$ such that $\tilde{f}|_{\mathbb{X}_{n}}=f$. We can extend $f$ by 0 outside $\mathbb{X}_{n}$: since $\mathrm{supp}(f)\subseteq\mathbb{X}_{n}$ is compact in $\mathbb{X}$, the extension by 0 is continuous and compactly supported because $\mathbb{X}_{n}$ is open in $\mathbb{X}$. Then for all $\xi\in M(\mathbb{X})$,
$\int_{\mathbb{X}_{n}}f\,d\,\pi_{n}(\xi)\;=\;\int_{\mathbb{X}_{n}}f\,d(\xi|_{\mathbb{X}_{n}})\;=\;\int_{\mathbb{X}}\tilde{f}\,d\xi.$
Since $\xi\mapsto\int_{\mathbb{X}}\tilde{f}\,d\xi$ is continuous on $(M(\mathbb{X}),\mathcal{T})$, it follows that $\xi\mapsto\int_{\mathbb{X}_{n}}f\,d\,\pi_{n}(\xi)$ is continuous, and hence $\pi_{n}$ is continuous.
Now let $A\in\bigcup_{n\geq 1}\phi_{n}^{-1}(\mathcal{T}_{n})$, so that $A=\phi_{n}^{-1}(A_{n})$ for some $A_{n}\in\mathcal{T}_{n}$. Using $\phi_{n}g=\pi_{n}$ and continuity of $\pi_{n}$ we obtain
$g^{-1}(A)=g^{-1}(\phi_{n}^{-1}(A_{n}))=(\phi_{n}g)^{-1}(A_{n})=\pi_{n}^{-1}(A_{n})\in\mathcal{T}.$
Therefore $g$ is continuous, and the proof is complete. ∎
@@
Let use denote by $\mathcal{M}_{n}$ the subspace $\sigma$-algebra of $M(\mathbb{X}_{n})$. Then by Theorem 2.2 any projective sequence $(\mu_{n})_{n\geq 1}$ of probability measures on $(M(\mathbb{X}_{n}),\mathcal{M}_{n})$ has a projective limit $\hat{\mu}_{\infty}$ which is a probability measure on the projective limit space $(M_{\mathbb{N}},\sigma(\mathcal{T}_{\mathbb{N}}))$, defined as in Theorem 2.1. By Proposition 5.1 this space is homeomorphic to $(M(\mathbb{X}),\mathcal{T})$ via the isomorphism $h$ and hence the probability measure $\hat{\mu}_{\infty}$ can be pushed forward to a probability measure $\mu_{\mathbb{N}}:=h\ast\hat{\mu}_{\infty}$ on $(M(\mathbb{X}),\sigma(\mathcal{T}))$, which equals $(M(\mathbb{X}),\mathcal{M})$, and hence $\mu_{\mathbb{N}}$ is a point process on $\mathbb{X}$.
@@
We need to show that the projective limit $\mu_{\mathbb{N}}$ is invariant with respect to the compatible direct pre-limit $\Phi_{\infty}$.
Recall that we have extended the action of $\Phi_{n}$ from $\mathbb{X}_{n}$ to $M(\mathbb{X}_{n})$ by defining $\varphi_{n}\xi_{n}(A)=\xi_{n}(\varphi_{n}^{-1}(A))$ and similarly for the action of $\Phi_{\infty}$ on $M(\mathbb{X})$.
Note that to prove invariance, it suffice to show that $\mu_{\mathbb{N}}$ is invariant with respect to $\eta_{m}\varphi_{m}$ for any $m\geq 1$ and $\varphi_{m}\in\Phi_{m}$. For this we follow a strategy similar to the proof of invariance in Theorem 3.1, making use of the fact that the projective limit measure $\mu_{\mathbb{N}}$ is the unique probability measure such that $\pi_{n}\ast\mu_{\mathbb{N}}=\mu_{n}$.
@@
Fix a $m\geq 1$, $\varphi_{m}\in\Phi_{m}$ and define $\hat{\mu}_{\infty}=\mu_{\mathbb{N}}\circ(\eta_{m}\varphi_{m})^{-1}$. Then we need to show that $\hat{\mu}_{\infty}=\mu_{\mathbb{N}}$, and by the uniqueness of projective limits it suffices to show that $\pi_{n}*\hat{\mu}_{\infty}=\mu_{n}$ holds for all $n\geq 1$. Let $k=\max\{m,n\}$ and $A_{n}\in\mathcal{M}_{n}$ be any measurable set of graphs with node labels in $\mathbb{X}_{n}$. Then,
@@
    where we used the commuting diagram in point 2 of the lemma for the third step.